OpenStudy (anonymous):

I need help with this Q! ^.^ What is the wavelength (in units of nm) of a blue light with a frequency of 6.41 x 10^14 Hz?

OpenStudy (anonymous):

@ganeshie8 :D

ganeshie8 (ganeshie8):

blue light is also light at what speed does the light travel ?

OpenStudy (anonymous):

Oh so it doesn't matter what color light it is? They all have the same speed?

ganeshie8 (ganeshie8):

Yes, all electromagnetic radiation travels at the same speed. Visible light is just a small part in the huge electromagnetic spectrum

ganeshie8 (ganeshie8):

However they will have different wavelengths and frequencies

ganeshie8 (ganeshie8):

so do you remember the speed of light ?

OpenStudy (anonymous):

Ahhh I see... so I have to use the formula \[\lambda =\frac{ v }{ f }\]

ganeshie8 (ganeshie8):

Yes, I am just trying to make you derive that formula. It follows trivially from the formula `distance = speed * time`

OpenStudy (anonymous):

Oh it does? :O

ganeshie8 (ganeshie8):

\(\lambda\) = wave length = `distance` of one cycle \(v\) = `speed` \(T\) = `time` taken to pass one cycle

ganeshie8 (ganeshie8):

distance = speed * time \(\lambda = v * T\)

ganeshie8 (ganeshie8):

recall the definition of frequency : number of cycles in one second therefore \(T = \dfrac{1}{f}\)

OpenStudy (anonymous):

Oh then my formula was wrong... because i divided them

OpenStudy (anonymous):

oh actually i take that back XD haha

OpenStudy (anonymous):

that makes sense!

ganeshie8 (ganeshie8):

your formula is correct too :) you're using ferquency, i have used timeperiod

OpenStudy (anonymous):

Aha! Right! lol :D

ganeshie8 (ganeshie8):

btw, normally we use \(c\) for the speed of light, not \(v\)

OpenStudy (anonymous):

so I use c= 3*10^8 ?

ganeshie8 (ganeshie8):

Yes

ganeshie8 (ganeshie8):

\[\lambda =\frac{ c }{ f } = \dfrac{3*10^8}{6.41*10^{14}} =?\]

OpenStudy (anonymous):

hmmm... it would be 3/6.41 * 10^-6 ?

OpenStudy (anonymous):

Is that exponent correct?

ganeshie8 (ganeshie8):

feed it to wolfram

OpenStudy (anonymous):

Awesomeness! Thanks so much!!

OpenStudy (anonymous):

A little math question!

ganeshie8 (ganeshie8):

do you know how to convert to nanometers ?

OpenStudy (anonymous):

I know it has something to with 10^-9

ganeshie8 (ganeshie8):

Yes, \(1nm=10^{-9}m\) so we get : \(4.68 * 10^{-7}m = 468*10^{-9}m = 468 nm\)

OpenStudy (anonymous):

how did you get 10^-7 again?

ganeshie8 (ganeshie8):

wolfram gave me that

OpenStudy (anonymous):

Ohhh i just got confused for a moment! Right right!! XD sorry haha

ganeshie8 (ganeshie8):

np, thats just math haha! it seems our calculation is right, wavelength of blue light is indeed around 468nm :